Nitrogen gas reacts with hydrogen gas to produce ammonia.
N2(g)+3H2(g)→2NH3(g)
How many grams of H2 are needed to react with 3.40 g of N2?

Thanks for your help!!
1
Expert's answer
2015-10-20T03:51:34-0400
The balanced chemical eqiation is
N2(g)+3H2(g)→2NH3(g)
The mass of nitrogen is m(N2) = 3.40 g
The amount of nitrogen is n(N2) = m(N2)/M(N2[]/sub) = 3.4/28.013 = 0.121 mol
The reacion stechiometry shows that we need 3 times more moles of hydrogen to react completely with this amount of nitrogen, or n(H[sub]2
) = 3x0.121 = 0.363 mol
The mass of hydrogen reqired is m(H2) = n(H2)M(H2) = 0.363x2.016 = 0.732 g
Answer: 0.732 g